OpenStudy (unicornpoopcookies):

is -1+9t^2 a difference of squares ?

OpenStudy (alexandervonhumboldt2):

-1+9t^2=9t^2-1 a^2-b^2=(a-b)*(a+b) 9t^2-1=(3^2)(t^2)-1^2

OpenStudy (alexandervonhumboldt2):

thus it is.......?

OpenStudy (alexandervonhumboldt2):

in your case a=9t^2 and b=1

OpenStudy (unicornpoopcookies):

It just asks for a yes or no answers

OpenStudy (alexandervonhumboldt2):

what do you think?

OpenStudy (unicornpoopcookies):

no.

jimthompson5910 (jim_thompson5910):

swap the `-1` and `9t^2` to get `9t^2 - 1` the 9t^2 can be written as (3t)^2

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):